OpenStudy (anonymous):

how do you type log base 6 in a calculator?

OpenStudy (chaise):

You can't - you have to convert to base 10 by using the formula. Do you need help with this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

log_6 (x) = log_10 (x)/log_10 (6)

OpenStudy (anonymous):

More generally log_a (x) = log_b (x)/log_b (a)

OpenStudy (anonymous):

what godess said. suppose you want \[\log_6(20)\] i.e. you want to solve \[6^x=20\] jfor x you would use \[\frac{\log(20)}{\log(6)}\]

OpenStudy (anonymous):

how you type it in depends on your calculator